Help
Sign up

Status

Check our services status at a glance

Title http7 hardware replacement
ID Operation #20
State completed
Beginning date 01/22/2012 midnight
End date 01/22/2012 12:20 a.m.
Concerned servers
  • http7

Messages

01/21/2012 3:55 p.m.

http7 disk cables and /dev/sda will be replaced tonight. The server will be unavailable during the operation.

01/22/2012 12:03 a.m.

The operation has started.

01/22/2012 12:09 a.m.

The operation completed successfully.

01/26/2012 11:36 a.m.

The underlying issue (complete disks lockup despite RAID1) is still present. /dev/sdb will be replaced tonight from 00:00 (CET). Expected downtime: less than 10 minutes.

01/27/2012 12:07 a.m.

The operation has started.

01/27/2012 12:11 a.m.

The operation completed successfully. Total downtime: 5 minutes.